if a function f is not defined at x=a that the limit
1
Expert's answer
2020-11-11T12:19:39-0500

The concept of limits has to do with the behaviour of the function close to x = a and not at x = a.

"\lim\limits_{x\to a}f(x)" may exist even if function f is undefined at "x=a"  


"f(x)=x, x\not=0"

"\lim\limits_{x\to 0}f(x)=0," "f" is undefined at "x=0."



"\lim\limits_{x\to a}f(x)" may be equal infinity even if function f is undefined at "x=a"  


"f(x)=\dfrac{1}{|x|},"

"\lim\limits_{x\to 0}f(x)=\infin," "f" is undefined at "x=0."


"\lim\limits_{x\to a}f(x)" may not exist if function "f" is undefined at "x=a"  


"f(x)=\dfrac{1}{x},"

"\lim\limits_{x\to 0^-}f(x)=-\infin, \lim\limits_{x\to 0^+}f(x)=\infin,"

"\lim\limits_{x\to 0}f(x)=\text{does not exist }"